OpenStudy (anonymous):

Log 4^9 My answer currently is Log 4^9 = 1.73205... Is it Correct ?

OpenStudy (anonymous):

9* Log 4 Use calculator now.

OpenStudy (anonymous):

5.4185....

OpenStudy (anonymous):

og 4^9 = 9*log 4 = 9*1.386294361119890618834464 = 12.476649250079015 here the base is 10

OpenStudy (anonymous):

@aron west log4 is not 1.3 its just .3

OpenStudy (anonymous):

I am sorry the base was natural logarithm.

OpenStudy (anonymous):

\[\log _{4}^{9}?\]